# How PARTs assemble into wholes: Learning the relative composition of images

### v1

@@
Abstract
The composition of objects and their parts, along with object-object positional relationships, provides a rich source of information for representation learning. Hence, spatial-aware pretext tasks have been actively explored in self-supervised learning. Existing works commonly start from a grid structure, where the goal of the pretext task involves predicting the absolute position index of patches within a fixed grid. However, grid-based approaches fall short of capturing the fluid and continuous nature of real-world object compositions. We introduce PART, a self-supervised learning approach that leverages continuous relative transformations between off-grid patches to overcome these limitations. By modeling how parts relate to each other in a continuous space, PART learns the relative composition of images—an off-grid structural relative positioning process that generalizes beyond occlusions and deformations. In tasks requiring precise spatial understanding such as object detection and time series prediction, PART outperforms strong grid-based methods like MAE and DropPos, while also maintaining competitive performance on global classification tasks with minimal hyperparameter tuning. By breaking free from grid constraints, PART opens up an exciting new trajectory for universal self-supervised pretraining across diverse datatypes—from natural images to EEG signals—with promising potential in video, medical imaging, and audio.
keywords:
@@ Abstract
We introduce PART, a self-supervised learning objective that predicts relative translations between random off-grid sampled patches.
We introduce a cross-attention relative encoder, that shares information across patches and imposes additional masking by predicting the relative translation for a subset of patch pairs.
The relative off-grid sampling introduces new capabilities such as off-grid reconstruction, extension to other aspect ratios and scales, patch uncertainty, and symmetry.
PART improves the downstream performance on precise tasks (i.e., localization and time-series prediction) while remaining competitive in classification.
@@ 1 Introduction
Most visual datasets lack the ground truth labels required for supervised learning [1, 2, 3, 4, 5]. However, even without relying on expensive labeled data, raw images provide a rich source of information for learning visual representations. Self-supervised learning leverages this information by defining pretext tasks [6]. While many pretext tasks focus on global visual invariances to pretrain deep networks [7, 8, 9, 10, 11, 12, 13, 14, 15], local spatial structures in images are suitable for precise downstream tasks [16]. Specifically, many works in self-supervised visual learning define local pretext tasks that extract patches from images using a grid structure, building on the popularity of grid-based vision transformers. For instance, early well-known self-supervised methods, often referred to as puzzle solvers [6, 17, 18] shuffle grid-shaped patches from the image, and the pretext task is to predict the absolute position of each patch, as in Jigsaw [19]. The same idea has been applied to pretraining the recent transformer architectures. For example, Masked Auto Encoders [20], MP3 [21], and DropPos [22] mask and shuffle grid-based patches, and regenerate the original unmasked image.
@@
So far, local self-supervised learning approaches, such as puzzle solvers and masked image modeling, (i) rely on fixed grid structures to patchify images and (ii) focus on absolute patch position prediction. However, real-world objects do not naturally conform to these rigid grid patterns, and the absolute position prediction limits the generalizability. In this work, we look beyond these two aspects that current self-supervised learning approaches adopt. First, we propose sampling patches freely, both in terms of location and size. By random off-grid sampling, we gain the ability to learn more fine-grained and occlusion-friendly representations. Second, we aim to learn the relative relationships between off-grid patches rather than the absolute relations of patches to the image.
@@
Such a relative pretext task allows us to learn visual representations better suited for downstream tasks beyond classification. Pretraining with absolute positions captures only the typical location of a given object in an image, leading to misinterpretations when objects appear in uncommon locations. Pretraining with relative positions enables the model to capture relationships between different objects and the internal composition of object parts, which can be generalized to other objects. Furthermore, while grid-based methods first sample the entire image uniformly and then apply a separate masking step, off-grid sampling inherently results in partial coverage—some regions are sampled while others are left out—thereby integrating masking directly into the sampling process. Off-grid sampling also allows greater control over the location and size of patches—for example, enabling focused sampling on objects, backgrounds, or uniformly across the image. Finally, off-grid sampling is more flexible and can be extended to patches of different scales and aspect ratios to also include odd-shaped objects. Figure 1 shows the difference between our approach and canonical grid-based self-supervised methods.
@@
We introduce PAirwise Relative Translations a pretraining method that predicts relative translations between randomly sampled patches. The pretext objective is set up as a regression task to predict the translation between each pair of patches (Figure 2). We also introduce a cross-attention architecture that serves as a projection head. We empirically show that PART outperforms baselines in precise tasks such as object detection and 1D EEG signal classification and remains competitive in image classification. We also perform ablation studies on different sampling strategies, projection head architectures, and the number of patch pairs. Code is openly available here 222https://anonymous.4open.science/r/predicting-relative-transforms-D1C0/README.md.
@@ 2 Related Work
Self-supervised learning (SSL) has emerged as a powerful paradigm for learning meaningful representations from vast amounts of unlabelled visual data, addressing the limitations posed by the scarcity of annotated datasets required for supervised learning [6, 23, 1, 2, 5]. By designing pretext tasks, SSL methods exploit the inherent structure of unlabelled images to learn discriminative features without human-annotated labels. These pretext tasks are generally divided into two broad categories: those that operate on the global structure of images and those that focus on local structures within the image. Global methods emphasize invariance to transformations applied to the entire image, while local methods focus on capturing fine-grained spatial structures.
@@
Global pretext tasks have been extensively explored in the literature and focus on leveraging image-wide transformations to learn invariant features. Notable global methods include learning to be invariant to rotations [7, 8, 9] and geometric transformations [24, 10, 11, 12]. In CIM [25], the model predicts the geometric transformation of a sample with respect to the original image. Other global tasks include colorization [13, 14], which aims to restore the original colors of grayscale images, and denoising [26, 27], where models are trained to predict and undo noise degradation. Another prominent task is instance discrimination [15, 28], where the goal is to distinguish between different objects. Building on instance discrimination, contrastive self-supervised methods have emerged, where different views or representations of the same data point are presented to one or two parallel models, with the objective of maximizing agreement between the two views. Common examples of this approach include MoCo [29, 30, 31] and SimCLR [32, 33] variants. A key concept in contrastive methods is the contrastive loss, which minimizes when the two input images are similar and maximizes when they are dissimilar [34, 35, 36]. The fundamental loss function enabling contrastive training for image-based SSL is InfoNCE [37, 38]. Self-labeling through clustering is another self-supervised approach, with methods such as DeepCluster [39], Self-Classifier [40], CoKe [41] and SwAV [42] employing this strategy. Distillation-based methods avoid the need for negative samples in clustering and contrastive techniques by training a student network to predict the representations of a teacher network [43, 44, 45, 46, 47]. Additionally, information-maximization methods focus on maximizing the information conveyed by decorrelated embeddings, eliminating the need for negative samples or asymmetric architectures [48, 49, 50, 51]. While these global methods effectively capture high-level patterns, they often miss the rich spatial details present in smaller image regions, which are critical for tasks involving fine-grained image understanding [16]. By focusing on transformations applied to the entire image, they may overlook the nuanced information embedded in local structures.
@@
Real-world images contain a wealth of local patterns, which are essential for understanding the finer aspects of visual content. Local pretext tasks, therefore, shift the focus towards modeling the internal spatial structure of images by extracting small grid-shaped patches from the image and defining pretext tasks on them. An early and prominent SSL method for learning local intra-image structures is the Jigsaw puzzle [17, 18], where patches are shuffled, and the model must predict their correct arrangement, commonly referred to as puzzle solvers. Masked image modeling (MIM) emerged as an adaptation of masked language modeling in NLP [52] to computer vision, introducing a new pretext task for self-supervised training [53]. In these methods, part of the input is masked, and the model is tasked with either reconstructing the original input or predicting the masked-out portion. This can be considered a variant of image inpainting. In Pathak et al. [54], the network is trained to inpaint the contents of a masked image region by understanding the context of the entire image. This approach has also been used to pretrain vision transformers [55], showing improved performance on downstream tasks compared to supervised and contrastive learning baselines.
@@
A popular example of masked image modeling is the MAE framework [20], based on BEiT [56], where a random subset of image patches is masked, and the pretext task involves reconstructing the entire image in pixel space. Similarly, in I-JEPA [57], the pretext task is to predict the representation of the rest of the image blocks given a single context block, focusing on large-scale models in linear classification. These methods can be categorized as generative-based approaches, where the model reconstructs the original input using generative models, such as VAEs [58] or GANs [59].
@@
However, generative-based masked prediction presents challenges, such as longer training times and the increased complexity of the reconstruction task [21]. To address these issues, alternative models have emerged that focus on predicting the absolute position of the masked patches instead of content reconstruction [21, 22]. For example, in MP3 [21] a random set of patches is masked with all the position embeddings, while in DropPos [22], the position embeddings of a random portion of the image are masked. In both methods, the pretext task is predicting the exact position of each patch, effectively solving the puzzle of determining where each patch originated from. This concept originates from earlier works [60, 61, 19, 62]. DILEMMA [63] enforces the prediction of positions for artificially misplaced patches, while Caron et al. [64] focuses on absolute position prediction for a random image portion, using the input image as a reference. In MP3 [21], the corresponding keys to a random set of patches are masked out, whereas, in DropPos [22], the position embeddings of a random portion of the image are masked out. The pretext task in both methods is predicting the exact position of each patch, requiring it to solve the puzzle of determining where each patch originated from. The idea behind these methods originates from works of Doersch et al. and Mundhenk et al. [60, 61] and later on the Jigsaw works [19, 62], where masking is performed by making a puzzle from a part of the image and pretraining a CNN to solve the jigsaw puzzle by predicting the absolute position of each piece. DILEMMA [63] enforces predicting the position of patches that have been artificially misplaced. In Caron et al. [64], the pretext task is the absolute position prediction of a random portion of the image given the input image as a reference. Self-supervised learning has demonstrated significant success across various domains and applications. After the introduction of Vision Transformers (ViT) [55] in 2021, ViTs were quickly adopted for self-supervised learning through methods like BEiT [56], MAE [20], and DINO [45], sparking considerable interest in leveraging these architectures for large-scale unlabeled data in self-supervised learning. While vision transformers often exhibit insensitivity to the order of input tokens [65, 66, 21, 22], suggesting that they tend to model relationships between unordered tokens, the aforementioned models focus explicitly on absolute grid-based position prediction. In contrast, PART is trained on predicting off-grid relative translations between random input patches.
@@
The notion of relative information has been applied in self-supervised learning across various tasks and domains. In graph representation learning, Peng et al. [67] proposed predicting the local relative contextual position of one node to another. For single-image depth estimation, Jiang et al. [68] introduced a method for estimating relative depth using motion from video sequences. In the domain of object detection, LIO [69] proposed a self-supervised spatial context learning module that captures the internal structure of objects by predicting the relative positions within the object’s extent. LIO localizes the main object in an image and learns the relative positions of other context points with respect to that main object. It operates by defining one main object as a reference, with other points or pixels learning their relative positions to this reference point. In contrast, PART learns relative information by considering any pair of patches as reference and target, allowing it to model relationships between any two patches.
3 Method: Pairwise Relative Translations
3.1 Random off-grid sampling
An overview of our method is shown in Figure 3. Given an image , we extract random patches from the image. and are the height and width of the image, and is the number of channels. With as the coordinates of the top left corner of the patch and as the coordinates of the bottom right corner of the patch, respectively. These patches are of shape and are in random positions of the image. Each sampled patch of shape is then resized to with channels. is the patch size, and is the number of patches. Now we have samples of that can be reshaped into the original image size . This reshaped image would be akin to a puzzled version of the original image if the random samples were on-grid and with a shape. In the process of off-grid random sampling, parts of the image are naturally masked out. In addition, some information about each patch’s spatial frequency is masked by resizing all samples to the patch size. The pretext task is set up such that the ViT [55], which processes images as sequences of patches, consumes images with incomplete information.
@@ 3.2 Relative translation parameterization
with and capturing relative position. In simple terms, the goal is to move the reference frame so that it translates into the target frame. In this context, when referring to a “frame”, we specify the bounding box itself rather than the actual pixel contents in the bounding box i.e. the patch (Figure 2). Note that our method differs from augmentations, which are applied to the entire image (e.g., rotations, scaling) to teach global invariance. In contrast, PART focuses on learning spatial relationships within the image, capturing the relative geometry between regions rather than enforcing global invariances.
@@
The emphasis on predicting the relative translation is crucial because information about the pixel space is lost after resizing to a uniform patch size. Here, the model no longer possesses details about the original image space. Thus, the two terms we seek to predict are the translation in normalized by the width of the reference patch and the translation in normalized by the height of the reference patch . In this case, both and are equal to patch size due to resizing.
3.3 Cross-attention relative encoder
The ViT model outputs a per-patch embedding . The relative encoder maps the per-patch embeddings to the relative translations between a random number of patch pairs (), resulting in . The two outputs per patch pair are the relative positions between the reference and target patches. Given , this module selects random index pairs () of patches with as the index of the reference patch and as the index of the target patch. The embeddings of reference patches and are then concatenated:
goes through a linear projection to convert from to . is fed into a cross-attention module [70] as the query, and is fed as both the key and the value. is the dimensionality of the keys/queries.
@@ 3.3 Cross-attention relative encoder
The cross-attention module allows for information dissemination between all patch embeddings and enables the model to focus on predicting the relative translation only for a subset of patch pairs. This imposes further masking of information given to the model. is only calculated for the subset of patch pairs. The model is trained with a mean squared error loss between the predicted relative translations and the ground-truth relative translations :
3.4 Supervised finetuning
@@
Once our model is pretrained, we can tune the network end-to-end using labeled data in a supervised setup. Following the standard ViT configuration, we eliminate the relative encoder and substitute it with a linear classification or detection head, after the [CLS] token. The [CLS] token aggregates global information about the input, as in the standard ViT. Unlike the pretraining phase, where no positional embedding is trained, we incorporate randomly initialized learnable position embeddings into the patch embeddings in this stage. Additionally, instead of the random sampling and masking in the pretraining phase, we perform fixed grid sampling when finetuning.
4 Implementation details
For both CIFAR-100 and ImageNet-1k, our pretraining and finetuning configurations closely align with the pipeline outlined in MP3 [21], which itself builds upon the foundational work of [71]. We adopt the MP3 codebase as our starting point for implementation. During the finetuning phase, we adhere strictly to the supervised training protocols recommended in DeiT [71], ensuring consistency with established practices. Detailed descriptions of the implementation specifics for each task are provided below:
Object detection
We evaluate the transfer learning capacity of our method on the COCO Dataset. We perform self-supervised pretraining on the ImageNet-1K [72] with a resolution of 224×224 using ViT-B [71] as the backbone. The model is pretrained for epochs with a learning rate of , a batch size of with patch pairs on GPUs. We perform end-to-end finetuning on COCO [73] for object detection. Specifically, Mask R-CNN [74] is finetuned with 1 schedule (12 epochs) and 10241024 resolution. We use the configuration of ViTDet [75] and take ViTB/16 [55] as the backbone. APb is reported as the main performance metric for object detection. We additionally report AP and AP.
1D time series classification
PhysioNet 2018 ”You Snooze You Win” Challenge Dataset [76] contains multi-channel biosignals that are continuously monitored overnight during sleep studies conducted at Massachusetts General Hospital. The task is to predict one of five sleep stages (Wake, Non-REM1, Non-REM2, Non-REM3, REM) given a 30-second window of data containing six channels of scalp electroencephalography (EEG). The EEG data is bandpass filtered with cutoffs 0.1-30 Hz and then re-sampled to have a 100 Hz sampling rate. Thirty-second windows are instance normalized as an additional pre-processing step before being tokenized by a linear layer. Forty 1-second patches are then randomly sampled. We use recordings from 1,653 subjects across the entire dataset for all pretraining strategies. For finetuning, we down-sample the dataset to 10 subjects to simulate a low-labeled data regime. For testing, the finetuned models are evaluated on recordings from 200 subjects, which are held out from both pretraining and finetuning stages. We randomly mask out 20% of the position embeddings for this experiment. All experiments use a 1D ViT backbone with 12.9M parameters and the input to the model is the 1D data for both pretraining and finetuning.
Image classification
We perform self-supervised pretraining on CIFAR100 [77] and ImageNet-1K [72] with a resolution of 32x32 and 224×224 respectively. Following Zhai et al. [21], we use ViT-S as the backbone of CIFAR100 and ViT-B [71] as the backbone of ImageNet-1K. During pretraining, we perform a hyperparameter search on learning rates , and the number of pairs and choose the best result for each experiment. We perform epochs of finetuning on CIFAR100 and epochs of finetuning on ImageNet-1K with a learning rate of . We report accuracy, error, and the mean squared error in and dimensions.
Choice of Architecture

### v2

@@
How PARTs assemble into wholes:
Learning the relative composition of images
1 Abstract
The composition of objects and their parts, along with object-object positional relationships, provides a rich source of information for representation learning. Hence, spatial-aware pretext tasks have been actively explored in self-supervised learning. Existing works commonly start from a grid structure, where the goal of the pretext task involves predicting the absolute position index of patches within a fixed grid. However, grid-based approaches fall short of capturing the fluid and continuous nature of real-world object compositions. We introduce PART, a self-supervised learning approach that leverages continuous relative transformations between off-grid patches to overcome these limitations. By modeling how parts relate to each other in a continuous space, PART learns the relative composition of images—an off-grid structural relative positioning that is less tied to absolute appearance and can remain coherent under variations such as partial visibility or stylistic changes. In tasks requiring precise spatial understanding such as object detection and time series prediction, PART outperforms grid-based methods like MAE and DropPos, while maintaining competitive performance on global classification tasks. By breaking free from grid constraints, PART opens up a new trajectory for universal self-supervised pretraining across diverse datatypes—from images to EEG signals—with potential in medical imaging, video, and audio.
@@ 2 Introduction
Most visual datasets lack the ground truth labels required for supervised learning [van2018inaturalist, zhou2018brief, 10.1001/jama.2017.14585, miller2016multimodal, 10.1145/3583138]. However, even without relying on expensive labeled data, raw images provide a rich source of information for learning visual representations. Self-supervised learning (SSL) leverages this information by defining pretext tasks [ozbulak2023know]. While many pretext tasks focus on global visual invariances to pretrain deep networks [rotnet, feng2019self, jing2018self, torpey2024affine, karimijafarbigloo2023self, musallam2024self, decolor, larsson2017colorization, wu2018unsupervised], local spatial structures in images are suitable for precise downstream tasks [park2023self]. Specifically, many works in self-supervised visual learning define local pretext tasks that extract patches from images using a grid structure, building on the popularity of grid-based vision transformers. For instance, early well-known SSL methods, often referred to as puzzle solvers [ozbulak2023know, carlucci2019domain, li2021jigsawgan] shuffle grid-shaped patches from the image and predict the absolute position of each patch, as in Jigsaw [jigsaw]. The same idea has been applied to pretraining the recent transformer architectures. For example, Masked Auto Encoders [mae], MP3 [mp3], and DropPos [droppos] mask and shuffle grid-based patches, and regenerate the original unmasked image.
@@
So far, local SSL approaches, such as puzzle solvers and masked image modeling, (i) rely on fixed grid structures to patchify images and (ii) predict absolute patch positions. Yet real-world objects rarely conform to rigid grid patterns, and absolute position prediction limits the generalizability. In this work, we look beyond these two aspects that current SSL approaches adopt. First, we propose sampling patches freely in location and size. The random off-grid sampling enables fine-grained and occlusion-friendly representations. Second, we learn the relative relationships between off-grid patches rather than the absolute relations of patches to the image.
@@
Such a relative pretext task allows us to learn visual representations better suited for downstream tasks beyond classification. Pretraining with absolute positions captures only the typical location of a given object in an image, leading to misinterpretations when objects appear in uncommon locations. Pretraining with relative positions enables the model to capture relationships between different objects and the internal composition of object parts, which can be generalized to other objects. Furthermore, while grid-based methods first sample the entire image uniformly and then apply a separate masking step, off-grid sampling inherently results in partial coverage—some regions are sampled while others are left out—thereby integrating masking directly into the sampling process. Off-grid sampling also allows greater control over the location and size of patches—for example, enabling focused sampling on objects, backgrounds, or uniformly across the image. Finally, off-grid sampling is more flexible and can be extended to patches of different scales and aspect ratios to also include odd-shaped objects. Figure 1 shows the difference between our approach and canonical grid-based SSL methods.
@@
We introduce PAirwise Relative Translations a pretraining method that predicts relative translations between randomly sampled patches. The pretext objective is set up as a regression task to predict the translation between each pair of patches (Figure 2). We also introduce a cross-attention architecture that serves as a projection head. We empirically show that PART outperforms baselines in precise tasks such as object detection and 1D EEG signal classification and remains competitive in image classification. We also perform ablation studies on different sampling strategies, projection head architectures, and the number of patch pairs. Code is openly available at https://github.com/Melika-Ayoughi/PART.
@@ 3 Related Work
Self-supervised learning has emerged as a powerful paradigm for learning meaningful representations from vast amounts of unlabelled visual data, addressing the limitations posed by the scarcity of annotated datasets required for supervised learning [ozbulak2023know, gui2024survey, van2018inaturalist, zhou2018brief, 10.1145/3583138]. By designing pretext tasks, SSL methods exploit the inherent structure of unlabelled images to learn discriminative features without human-annotated labels. These pretext tasks are generally divided into two broad categories: those that operate on the global structure of images and those that focus on local structures within the image. Global methods emphasize invariance to transformations applied to the entire image, while local methods focus on capturing fine-grained spatial structures.
@@
Global pretext tasks have been extensively explored in the literature and focus on leveraging image-wide transformations to learn invariant features. Notable global methods include invariance to rotations [rotnet, feng2019self, jing2018self] and geometric transformations [zhang2019aet, torpey2024affine, karimijafarbigloo2023self, musallam2024self]. In CIM [li2023correlational], the model predicts the geometric transformation of a sample with respect to the original image. Other global tasks include colorization [decolor, larsson2017colorization], denoising [noise, yasarla2024self], and instance discrimination [wu2018unsupervised, zhao2020makes]. Building on instance discrimination, contrastive self-supervised methods have emerged, where different views or representations of the same data point are presented to one or two parallel models, with the objective of maximizing agreement between the two views. Common examples include MoCo [moco, chen2020improved, chen2021empirical] and SimCLR [simclr, simclr2] variants. A key concept in contrastive methods is the contrastive loss, which minimizes when the two input images are similar and maximizes when they are dissimilar [chopra2005learning, hadsell2006dimensionality, zhang2022dual]. The fundamental loss function enabling contrastive training for image-based SSL is InfoNCE [sohn2016improved, oord2018representation]. Self-labeling through clustering is another SSL approach, with methods such as DeepCluster [caron2018deep], Self-Classifier [amrani2022self], CoKe [qian2022unsupervised] and SwAV [swav]. Distillation-based methods avoid the need for negative samples in clustering and contrastive techniques by training a student network to predict the representations of a teacher network [cpc, byol, dino, chen2021exploring, ren2023tinymim]. Additionally, information-maximization methods focus on maximizing the information conveyed by decorrelated embeddings, eliminating the need for negative samples or asymmetric architectures [bardes2022vicregl, zbontar2021barlow, kalantidis2021tldr, zhang2022align]. While these global methods effectively capture high-level patterns, they often miss the rich spatial details present in smaller image regions, which are critical for tasks involving fine-grained image understanding [park2023self]. By focusing on transformations applied to the entire image, they may overlook the nuanced information embedded in local structures.
@@
Real-world images contain a wealth of local patterns, which are essential for understanding the finer aspects of visual content. Local pretext tasks, therefore, shift the focus towards modeling the internal spatial structure of images by extracting small grid-shaped patches from the image and defining pretext tasks on them. An early SSL method for learning local intra-image structures is the Jigsaw puzzle [carlucci2019domain, li2021jigsawgan], where patches are shuffled, and the model must predict their correct arrangement, commonly referred to as puzzle solvers. Masked image modeling (MIM) emerged as an adaptation of masked language modeling in NLP [bert], introducing a new pretext task for self-supervised training [chen2020generative]. In these methods, part of the input is masked, and the model is tasked with either reconstructing the original input or predicting the masked-out portion. This can be considered a variant of image inpainting. In Pathak et al. [contextencoder], the network is trained to inpaint the contents of a masked image region by understanding the context of the entire image. This approach has also been used to pretrain vision transformers [vit], showing improved performance on downstream tasks compared to supervised and contrastive learning baselines.
@@
A popular example of masked image modeling is MAE [mae], based on BEiT [beit], where a random subset of image patches is masked, and the model reconstructs the entire image in pixel space. Similarly, in I-JEPA [ijepa], the pretext task is to predict the representation of the rest of the image blocks given a single context block, focusing on large-scale models in linear probing. These methods can be categorized as generative-based approaches, where the model reconstructs the original input using generative models, such as VAEs [vae] or GANs [goodfellow2020generative].
@@
However, generative-based masked prediction presents challenges, such as longer training times and increased complexity of the reconstruction task [mp3]. To address these issues, alternative models focus on predicting the absolute position of the masked patches instead of pixel reconstruction [mp3, droppos]. In MP3 [mp3], the corresponding keys to a random set of patches are masked out, whereas, in DropPos [droppos], the position embeddings of a random portion of the image are masked out. The pretext task in both methods is predicting the exact position of each patch, requiring it to solve the puzzle of determining where each patch originated from. The idea behind these methods originates from works of Doersch et al. and Mundhenk et al. [doersch2015unsupervised, contextprediction2] and later on the Jigsaw works [jigsaw, jigsaw++], where masking is performed by making a puzzle from a part of the image and pretraining a CNN to solve the jigsaw puzzle by predicting the absolute position of each piece. DILEMMA [sepehr] enforces predicting the position of patches that have been artificially misplaced. In Caron et al. [loca], the pretext task is the absolute position prediction of a random portion of the image given the input image as a reference. Self-supervised learning has demonstrated significant success across various domains and applications. After the introduction of Vision Transformers (ViT) [vit] in 2021, ViTs were quickly adopted for self-supervised learning through methods like BEiT [beit], MAE [mae], and DINO [dino], sparking considerable interest in leveraging these architectures for large-scale unlabeled data in self-supervised learning. While vision transformers often exhibit insensitivity to the order of input tokens [chenempirical, naseer2021intriguing, mp3, droppos], suggesting that they tend to model relationships between unordered tokens, the aforementioned models focus explicitly on absolute grid-based position prediction. In contrast, PART is trained on predicting off-grid relative translations between random input patches.
@@
The notion of relative information has been applied in self-supervised learning across various tasks and domains. In graph representation learning, Peng et al. [graph] proposed predicting the local relative contextual position of one node to another. For single-image depth estimation, Jiang et al. [video] introduced a method for estimating relative depth using motion from video sequences. In the domain of object detection, LIO [od] proposed a self-supervised spatial context learning module that captures the internal structure of objects by predicting the relative positions within the object’s extent. LIO localizes the main object in an image and learns the relative positions of other context points with respect to that main object. It operates by defining one main object as a reference, with other points or pixels learning their relative positions to this reference point. Similarly, HASSOD [cao2023hassod] identifies objects in an image through clustering and progressively refines object understanding hierarchically by discovering object parts. Both approaches are tailored to object-centric tasks. In contrast, PART is applicable to both object-centric and object-agnostic data, such as EEG, since it learns relative information by considering any pair of patches as reference and target, allowing it to model relationships between any two patches.
4 Method
Sampling
An overview of our method is shown in Figure 3. Given an image , we extract random patches from the image. and are the height and width of the image, and is the number of channels. With as the coordinates of the top left corner of the patch uniformly sampled across the image and as the coordinates of the bottom right corner of the patch, respectively. These patches are of shape and are in random positions of the image. is the patch size, and is the number of patches. Now we have samples of that can be reshaped into the original image size . This reshaped image would be akin to a puzzled version of the original image if the random samples were on-grid and with a shape. In the process of off-grid random sampling, parts of the image are naturally masked out. In addition, some information about each patch’s spatial frequency is masked by resizing all samples to the patch size. The pretext task is set up such that the ViT [vit], which processes images as sequences of patches, consumes images with incomplete information.
@@ 4 Method
[…]
@@
The emphasis on predicting the relative translation is key because the pixel space information is lost after resizing patches to a uniform size. After resizing, the model no longer possesses details about the original image space and needs to learn to be robust to different image resolutions. The two terms we seek to predict are the translation in normalized by the width of the reference patch and the translation in normalized by the height of the reference patch . In this case, both and are equal to patch size due to resizing.
Relative encoder architecture
@@
The ViT model outputs a per-patch embedding . The relative encoder maps the per-patch embeddings to the relative translations between a random number of patch pairs (), resulting in . The two outputs per patch pair are the relative positions between the reference and target patches. Given , this module selects random index pairs () of patches with as the index of the reference patch and as the index of the target patch. The embeddings of reference patches and are then concatenated: . goes through a linear projection to convert from to . is fed into a cross-attention module [attention] as the query, and is fed as both the key and the value. is the dimensionality of the keys/queries: . The cross-attention module allows for information dissemination between all patch embeddings and enables the model to focus on predicting the relative translation only for a subset of patch pairs. This imposes further masking of information given to the model. We discuss the pros and cons of this design choice in depth in the ablations section. is only calculated for the subset of patch pairs. The model is trained with a mean squared error loss between the predicted relative translations and the ground-truth relative translations : .
Training setup
@@
Once the model is pretrained, we tune the network end-to-end using labeled data in a supervised setup. Following the standard ViT setup, we eliminate the relative encoder and substitute it with a linear classification or detection head after the [CLS] token, which aggregates global input information. Unlike pretraining, we incorporate randomly initialized learnable position embeddings and apply fixed grid sampling instead of random sampling and masking.